OpenStudy (anonymous):

Flying to Kampala with a tailwind a plane averaged 158 km?h/ On the return trip the plane only averaged 112km/h while flying back into the same wind. Find the speed of the wind and the speed of the plane in still air. *needs two equations*

OpenStudy (vishal_kothari):

The averages are given. So that means the first trip is the equivalent of of 158 miles in one hour. Since this is with the wind, using the distance = rate x time, we have : 158 = (a + w) 1, where a = airspeed, w - wind speed, and 1 is the time. For the second one, 112 = (a - w) x 1, a = airspeed, etc. 158 = a + w 112 = a - w Add the two equations 270 = 2a 135 = a So the airspeed is 135 km/h 158 = 135 + w 23 = w The windspeed is 23 km/h